### v1

# Adversary Resilient Learned Bloom Filters

Abstract
Creating an adversary resilient Learned Bloom Filter [1] with provable guarantees is an open problem [2]. We define a strong adversarial model for the Learned Bloom Filter. We also construct two adversary resilient variants of the Learned Bloom Filter called the Uptown Bodega Filter and the Downtown Bodega Filter. Our adversarial model extends an existing adversarial model designed for the Classical (i.e not “Learned”) Bloom Filter by Naor Yogev [3] and considers computationally bounded adversaries that run in probabilistic polynomial time (PPT). We show that if pseudo-random permutations exist, then a secure Learned Bloom Filter may be constructed with $\lambda$ extra bits of memory and at most one extra pseudo-random permutation in the critical path. We further show that, if pseudo-random permutations exist, then a high utility Learned Bloom Filter may be constructed with $2\lambda$ extra bits of memory and at most one extra pseudo-random permutation in the critical path. Finally, we construct a hybrid adversarial model for the case where a fraction of the workload is chosen by an adversary. We show realistic scenarios where using the Downtown Bodega Filter gives better performance guarantees compared to alternative approaches in this hybrid model.
1 Introduction
The Bloom Filter is a probabilistic data structure that solves the Approximate Membership Query Problem. The data structure now known as the “Bloom” Filter was initially proposed as method $2$ in the section “Two Hash-Coding Methods with Allowable Errors” in a 1970 paper by Burton H. Bloom [4][5]. The Bloom Filter has applications in databases, cryptography, computer networking, social networking [6], and network security [7]. The Learned Bloom Filter is a novel data structure first proposed by Kraska et al [1] in 2017. The Learned Bloom Filter can be thought of as a Bloom Filter working in collaboration with a Learning Model. There are currently no known provably secure constructions of the Learned Bloom Filter. Prior work [2] has left the security of the Learned Bloom Filter as an open problem.
In this section, we first summarize the original formulation of the Bloom Filter and then provide background on the Learned Bloom Filter. Next, we outline the scope of this work and provide a summary of our results. Lastly, we provide motivation for our results and discuss related work.
1.1 The Bloom Filter
A Bloom Filter representing a set $S$ may have false positives ($s\notin S$ may return true) but does not have false negatives ($s\in S$ is always true). The Learned Bloom Filter provides the same correctness guarantees as the Bloom Filter, but with potentially better performance for the same memory budget. We first discuss the (Classical) Bloom Filter, and then discuss the Learned Bloom Filter.
1.1.1 Classical Bloom Filter
Figure 1 provides a helpful illustration of a Bloom Filter and the insert and check operations we introduce below.
Figure 1: Example of a Bloom Filter with $m$ = 8 and $k$ = 2. Initially, all $m$ bits are unset. Each element $x_{i}$ is hashed $k$ times, and each corresponding bit is set. To check each element $y_{i}$, the element is hashed $k$ times. If any corresponding bit is unset, the element $y_{i}$ is not in set $S$ (with probability $1$). If all corresponding bits are set, the element $y_{i}$ is either in set $S$ or the element $y_{i}$ has caused the Bloom Filter to return a false positive
Definition 1.1 (Bloom Filter).
A Bloom Filter for representing set $S$ with cardinality $n$ is a zero initialized array of $m$ bits. A Bloom Filter requires $k$ independent hash functions $h_{i}$ such that the range of each $h_{i}$ is the set of integers $\{1,\ldots,m\}$ [7].
Most mathematical treatments such as Mitzenmacher and Broder [7] make the convenient assumption that each $h_{i}$ maps each item in the universe to a random number uniformly over the (integer) range $[1,m]$. In the remainder of this text, we shall refer to this formulation of the Bloom Filter as the Classical Bloom Filter. This is to distinguish it from the Learned Bloom Filter, which our results are focused on.
Operation 1 (Insert).
For each element $x\in S$, the bits $h_{i}(x)$ are set to $1$ for $i\in[1,k]$.
If a bit already set to $1$ is set to $1$ again, its value remains $1$ i.e a double set does not flip the bit back to 00.
Operation 2 (Check).
For an element $x$, we return true if all $h_{i}(x)$ map to bits that are set to $1$. If there exists an $h_{i}(x)$ that maps to a bit that is 00, we return false.
Figure 2: Example of a Learned Bloom filter with a Learned Oracle and a Backup Classical Bloom filter that only checks values that are with high probability negative in the Learned Oracle, to ensure a one-sided error bound (i.e only false positives and no false negatives).
1.1.2 Learned Bloom Filter
The Learned Bloom Filter is a novel data structure proposed by Kraska et al [1] in 2017. Kraska et al suggest using a pre-filter ahead of the Classical Bloom Filter, where the pre-filter is derived from a learning model. The learning model estimates the probability of an element $x$ being in the set $S$. This allows the use of a smaller (in terms of memory) Classical Bloom Filter compared to the case where a Classical Bloom Filter is used alone. A mathematical model and guarantees for the Learned Bloom Filter were first provided by Mitzenmacher [8] in 2018. While Kraska et al used a neural network for the pre-filter, Mitzenmacher generalizes the pre-filter to use any approach that estimates the probability of an element $x$ being in the set $S$. We use this generalized formulation of the Learned Bloom Filter in this work. We introduce the Learned Bloom Filter using a self-contained definition adapted from Mitzenmacher [8]. We extend this definition in Section 2.2.
Definition 1.2 (Learned Bloom Filter).
A Learned Bloom filter on a set of positive keys $\mathcal{K}$ and negative keys $\mathcal{U}$ is a function $f:U\mapsto[0,1]$ and threshold $\tau$ , where $U$ is the universe of possible query keys, and an associated Classical Bloom Filter $B$, referred to as a Backup Classical Bloom filter. The Backup Classical Bloom Filter holds the set of keys $\{z:z\in\mathcal{K},f(z)<\tau\}$. For a query $y$, the Learned Bloom filter returns that $y\in\mathcal{K}$ if $f(y)\geq\tau$ , or if $f(y)<\tau$ and the Backup Classical Bloom Filter returns that $y\in\mathcal{K}$. The Learned Bloom filter returns $y\notin\mathcal{K}$ otherwise.
The Learned Bloom Filter provides better performance on the false positive rate while maintaining the guarantee of having no false negatives. We show an example of a Learned Bloom Filter adapted from Mitzenmacher et al [8] in figure 2.
1.2 Our Results
We introduce an adversarial model for the Learned Bloom Filter. Our model extends the adversarial model for the Classical Bloom Filter introduced by Naor Yogev [3]. Using our adversarial model, we introduce the notion of an adversarial resilient Learned Bloom Filter, as well as an adversarial reveal resilient Learned Bloom Filter. The precise definitions and the adversarial model we introduce are provided in Section 2.
1.2.1 Secure Learned Bloom Filter
We introduce the first provably secure constructions of the Learned Bloom Filter and show that using pseudorandom permutations, one can create provably secure constructions of the Learned Bloom Filter with at most $2\lambda$ extra bits of memory, where $\lambda$ is the security parameter.
Theorem 1.3 (Secure Constructions).
Let $B$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, we show that
  1. 1.
    There exists a negligible function $\mathrm{negl}(\cdot)$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial resilient Learned Bloom Filter that uses $m^{\prime}=m+\lambda$ bits of memory
  2. 2.
    There exists a negligible function $\mathrm{negl}(\cdot)$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient Learned Bloom Filter that uses $m^{\prime}=m+2\lambda$ bits of memory
These results are formally proved in Theorem 3.7 and Theorem 3.10 of Section 3.
1.2.2 Utility of the Secure Learned Bloom Filter
To explore the utility of the Secure Learned Bloom Filter, we introduce a hybrid model where an adversary is allowed to choose $\alpha N$ queries out of a workload of $N$ queries sent to the Bloom Filter.
Theorem 1.4 (Performance in the Hybrid Model).
For a given memory budget $M=m_{L}+m_{A}+m_{B}+2\lambda$, any set $S$ and corresponding training dataset $\mathcal{D}$, in the hybrid adversarial setting with given $\alpha=\alpha_{P}+\alpha_{N}$, the expected false positive probability of a Secure Learned Bloom Filter is lower than the expected false positive probability of a Secure Classical Bloom Filter if the following holds true:
$\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)<\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, where the $FPR$ is a function for the false positive probability of a Secure Classical Bloom Filter, and $FPR_{DB}$ is the function for the false positive probability of a Secure Learned Bloom Filter.
8798These results are formally proven in Theorem 5.3 of Section 5. In this result, when we refer to a “Secure Learned Bloom Filter”, we refer in particular to the “Downtown Bodega Filter” construction we define in Section 3.3. We demonstrate multiple settings in a hybrid adversarial model where the Secure Learned Bloom Filter out-performs the Secure Classical Bloom Filter (Section 5.4). The advantage of the Secure Learned Bloom Filter compared to the Secure Classical Bloom Filter intricately depends on the number of adversarial queries and the performance of the Learned Oracle.
1.3 Motivation
The Bloom Filter and its variants have numerous applications in computing [9] [7]. We borrow discussion on Bloom Filter applications from the survey by Tarkoma et al [9]. The Bloom Filter may be implemented in kernel space in a Linux network driver for performant filtering of network packets. Loop detection in network protocols, and multicast forwarding engines may also utilize the Bloom Filter. Deep Packet Scanners and Packet Classifiers have also found the Bloom Filter helpful for improving efficiency. The Bloom Filter may be used to detect heavy flows in network traffic from the vantage point of a router. The Bloom Filter has also been used in the OPUS system [9] that stores a list of words that involve poor password choices encouraging users to select better passwords. The Bloom Filter has also found success in the detection of hash tampering in network attached disks. Google’s BigTable system uses the Bloom Filter to minimize disk reads. Apache Hadoop also uses the Bloom Filter as an optimization in the reduce stage of its map/reduce implementation. Other applications of the Bloom Filter include uses in the realms of peer-to-peer networking, and caching.
1.3.1 The Need for Security
A large number of the applications of the Bloom Filter involve critical infrastructure [10]. It is possible to forge false positives in a naively implemented Bloom Filter [10] allowing an adversary to make the Bloom Filter deviate from its behavior. Gerbet et al [10] show practical attacks on the Scrapy web-spider, the Bitly Dablooms spam filter, and the Squid web cache. Naor and Yogev [3] motivate the need for securing the Bloom Filter by considering a white-list of email addresses for the purposes of spam filtering. In their scenario, an adversary that can forge false positives may easily infiltrate the spam filter.
1.4 Related Work
Section 1.4.1 provides a thorough overview of prior work on adversarial models and security of the classical Bloom Filter. Similarly, Section 1.4.2 discusses prior work on the security of the Learned Bloom Filter.
1.4.1 Classical Bloom Filter
Gerbet et al [10] suggest practical attacks on the Classical Bloom Filter and the use of universal hash functions and message authentication codes (MACs) to mitigate a subset of those attacks. Naor and Yogev [3] define an adversarial model for the Classical Bloom Filter and use it to prove that $(1)$ for computationally bounded adversaries, non-trivial adversary resilient Bloom filters exist if and only if one-way functions exist, and $(2)$ for computationally unbounded adversaries, there exists a Classical Bloom Filter that is secure against $t$ queries while using only $\mathcal{O}(n\log\frac{1}{\epsilon}+t)$ bits of memory. $n$ is the size of the set and $\epsilon$ is the desired error. We borrow their idea of using Pseudorandom Permutations for the Classical Bloom Filter and apply it to the Learned Bloom Filter.
Clayton et al [11] analyze not only the Classical Bloom Filter, but also the Counting Bloom Filter, and the Count-Min Sketch, in an adversarial setting. Clayton et al use a stronger adversarial model than Naor and Yogev [3], allowing an adversary to perform insertions and giving an adversary access to the internal state of the Classical Bloom Filter. Clayton et al propose the use of salts and keyed pseudo-random functions for securing the Classical Bloom Filter. They do not address Learned Probabilistic Data Structures including the Learned Bloom Filter. Both Naor and Yogev, and Clayton et al, perform their analysis in a game-based setting.
Filic et al [12] investigate the adversarial correctness and privacy of the Classical Bloom Filter and an insertion-only variant of the Cuckoo Filter. They use a stronger adversarial model than Naor and Yogev [3] allowing an adversary to insert entries into the Classical Bloom Filter and query for the internal state of the Classical Bloom Filter. Unlike our work, Filic et al [12] do not address the adversarial correctness of the Learned Bloom Filter or its variants. Filic et al [12] perform their analysis in a simulator-based setting.
1.4.2 Learned Bloom Filter
The authors are only aware of one prior work that addresses the Learned Bloom Filter in an adversarial setting, Reviriego et al [2]. They propose a practical attack on the Learned Bloom Filter. They suggest two possible mitigations for their proposed attack: swapping to a Classical Bloom Filter upon detection of the attack, or adding a second Backup Classical Bloom Filter. However, they do not provide any provable guarantees on the performance of the Learned Bloom Filter in the presence of adversaries. They leave the security of the Learned Bloom Filter as an open problem in their work.
2 Adversarial Model
We first describe the adversarial model of Naor and Yogev [3] for the Classical Bloom Filter and then use it as the basis for creating an adversarial model for the Learned Bloom Filter. We refer to the adversarial model defined by Naor and Yogev [3] as the classical adversarial model. Section 2.1 contains a treatment of the classical adversarial model. We also introduce a stronger adversary than the one described in Naor and Yogev’s [3] model that has access to the internal state of the Classical Bloom Filter. Section 2.2 introduces a definition of the Learned Bloom Filter adapted from [8] and discusses extensions to the classical adversarial model to make it work with the Learned Bloom Filter.
2.1 Classical Adversarial Model
Let $S$ be a finite set of cardinality $n$ in a suitable finite universe $U$ of cardinality $u$. Let $M$ be a compressed representation of $S$. Let $r$ be any random string and $M^{S}_{r}$ be a compressed representation of $S$ with $r$. Let $\lambda$ be a security parameter. Let $A=(A_{C},A_{Q})$ be any probabilistic polynomial time (PPT) adversary.
Definition 2.1 (Construction).
We define $C$ to be a setup algorithm such that $C(1^{\lambda},S)=M$. We define $C_{r}$, the randomized version of $C$, to be a setup algorithm such that $C_{r}(1^{\lambda},S)=M^{S}_{r}$. Note that $M$ and $M_{r}^{S}$ are both compressed representations of $S$, as defined above.
As a running example, consider the set $S=\{x\}$ in a Classical Bloom Filter that uses $2$ hash functions, $h_{1},h_{2}$ and $4$ bits such that $h_{1}(x)=1$ and $h_{2}(x)=3$. A trivial deterministic setup algorithm, on input $S$, would then generate the representation $M^{S}_{r}=1010$.
Definition 2.2 (Query).
We are provided a set $S$ and a compressed representation of that set, $M_{r}^{S}$ (where $r$ is any random string). We define $Q_{s}$ to be a query algorithm such that $Q_{s}(M_{r}^{S},x)=1$ if $x\in S$, and $Q_{s}(M_{r}^{S},x)\in\{0,1\}$ if $x\notin S$. $Q_{s}$ may not be randomized and may not change $M_{r}^{S}$.
In our running example, a trivial query algorithm returns $1$ if and only if all hashes for an element return indexes that are set i.e $Q_{s}(M^{S}_{r},x)=(M_{r}^{S}(h_{1}(x))=1\land M_{r}^{S}[h_{2}(x)]=1)$. Consider a new element $y$ for which $h_{1}(y)=1$ and $h_{2}(y)=2$. With $M^{S}_{r}=1010$, $Q(M^{S}_{r},x)$ returns $1$ since both indices $1$ and $3$ are set, however $Q(M^{S}_{r},y)$ return 00 as index $2$ is not set.
Definition 2.3.
We define $Q_{u}$ to be a query algorithm similar to $Q_{s}$ differing only in that $Q_{u}$ may be randomized and it may change the compressed representation of $S$, $M_{r}^{S}$, after each query.
We now give a precise definition for the Classical Bloom Filter in an adversarial setting.
Definition 2.4 (Classical Bloom Filter).
Let a Classical Bloom Filter be a data structure $B=(C_{r},Q)$ where $C_{r}$ obeys Definition 2.1 and $Q$ obeys either Definition 2.2 or Definition 2.3.
We define a special class of Classical Bloom Filters which were coined “steady” Classical Bloom Filters by Naor and Yogev [3]. Steady Classical Bloom Filters do not change their internal representation $M^{S}_{r}$ after the setup algorithm $C_{r}$ has executed. In other words, only query algorithms of the type $Q_{s}$ are permitted in the steady setting and query algorithms of the type $Q_{u}$ are not permitted.
Definition 2.5 (Steady).
Let a steady $(n,\epsilon)$-Classical Bloom Filter be a Classical Bloom Filter $B_{s}=(C_{r},Q_{s})$ such that $Q_{s}$ obeys Definition 2.2 and $\forall x\in U$, it holds that
  1. 1.
    Completeness: $\forall x\in S:P[Q_{s}(C_{r}(S),x)=1]=1$
  2. 2.
    Soundness: $\forall x\notin S:P[Q_{s}(C_{r}(S),x)=1]\leq\epsilon$
where the probabilities are taken over $C_{r}$.
Referring again to our running example, our trivial query algorithm is complete as for element $x$ which already exists in $S$, both index $1$ ($h_{1}(x)=1$) and index $3$ ($h_{2}(x)=3$) are set, so our trivial query algorithm returns $1$. Our trivial query algorithm is also sound because for any $y\neq x$, $P[h_{1}(y)\in\{1,3\}]$ is bounded and $P[h_{2}(y)\in\{1,3\}]$ is bounded, therefore the probability of our trivial query algorithm returning $1$ is bounded.
Now we construct our first adversarial challenge for the Classical Bloom Filter in the steady setting. A probabilistic polynomial time (PPT) adversary $A$, as defined at the start of this section, is given a security parameter $1^{\lambda+n\log(u)}$ and is allowed to construct a set $S$. The set $S$ is then given to construction algorithm $C_{r}$ along with the security parameter to yield representation $M^{S}_{r}$. The adversary is allowed $t$ queries to the query algorithm $Q_{s}$ for which it is provided the results. After the $t$ queries, the adversary must output an element $x^{*}$. If $x^{*}$ is a false positive, and has not been queried before, the adversary wins the challenge. Otherwise, the adversary loses the challenge. We define this precisely in Challenge 1
Challenge 1 (Resilient).
We denote this challenge as ${\Lambda^{1}}_{A,t}(\lambda)$.
  1. 1.
    $S\leftarrow A_{C}(1^{\lambda+n\log(u)})$
  2. 2.
    $M^{S}_{r}\leftarrow C_{r}(1^{\lambda+n\log(u)},S)$
  3. 3.
    $x^{*}\leftarrow A_{Q}^{Q_{s}(M^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S)$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $Q_{s}(M^{S}_{r},\cdot)$.
  4. 4.
    If $x^{*}\notin S\cup\{x_{1},\ldots,x_{t}\}$ and $Q_{s}(M^{S}_{r},x^{*})=1$, output $1$. Otherwise, output 00.
We now define an adversarial resilient Classical Bloom Filter based on the random variable ${\Lambda^{1}}_{A,t}(\lambda)$.
Definition 2.6 (Resilient).
Let an $(n,t,\epsilon)$-adversarial resilient steady Classical Bloom Filter be any steady Classical Bloom Filter for which it holds that, $\forall\lambda>n\in\mathbb{N},P[\Lambda^{1}_{A,t}(\lambda)=1]\leq\epsilon$.
Now, we create an extension to Naor and Yogev’s [3] model, introducing a stronger adversary that has access to the internal state of the Classical Bloom Filter. We construct our second adversarial challenge for the Classical Bloom Filter in the steady setting. This challenge is almost identical to the first challenge with the only difference being that the adversary is allowed access to the representation $M_{r}^{S}$. We define our second challenge more precisely in Challenge 2.
Challenge 2 (Reveal Resilient).
We denote this challenge as ${\Lambda^{2}}_{A,t}(\lambda)$.
  1. 1.
    $S\leftarrow A_{C}(1^{\lambda+n\log(u)})$
  2. 2.
    $M^{S}_{r}\leftarrow C_{r}(1^{\lambda+n\log(u)},S)$
  3. 3.
    $A_{Q}$ is allowed access to $M_{r}^{S}$
  4. 4.
    $x^{*}\leftarrow A_{Q}^{Q_{s}(M^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S)$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $Q_{s}(M^{S}_{r},\cdot)$.
  5. 5.
    If $x^{*}\notin S\cup\{x_{1},\ldots,x_{t}\}$ and $Q_{s}(M^{S}_{r},x^{*})=1$, output $1$. Otherwise, output 00.
Analogous to Definition 2.6, we now define an adversarial reveal resilient Classical Bloom Filter based on the random variable $\Lambda^{2}_{A,t}(\lambda)$ (from Challenge 2).
Definition 2.7 (Reveal Resilient).
Let an $(n,t,\epsilon)$-adversarial reveal resilient steady Classical Bloom Filter be any steady Classical Bloom Filter for which it holds that, $\forall\lambda>n\in\mathbb{N},P[\Lambda^{2}_{A,t}(\lambda)=1]\leq\epsilon$.
2.2 Learned Adversarial Model
In this section, we first discuss a mathematical model for the Learned Bloom Filter. We then create an adversarial model based on the Learned Bloom Filter. We define challenges and security definitions for the Learned Bloom Filter that are analogous to the ones we defined for the Classical Bloom Filter in Section 2.1.
Mitzenmacher [8] was the first to create a mathematical model for the Learned Bloom Filter, in 2018. Our model is heavily based on Mitzenmacher’s model but with some additional definitions to suit our adversarial setting. Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a training dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$.
Definition 2.8 (Dataset Construction).
Let $\Delta_{r}$ be any construction algorithm that takes a set $S$, and constructs a training dataset $\mathcal{D}$ for $S$.
Definition 2.9 (Learning Model).
Let a Bloom Filter Learning Model, $l:U\mapsto[0,1]$, be any function that maps elements in a suitable finite universe to a probability.
We train a Bloom Filter Learning Model, $l_{r}$, on $\mathcal{D}$. Let $l_{r}(x)$ be the probability estimate from the learning model that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ on to the Backup Classical Bloom Filter. Figure 2 provides a helpful illustration. In many places in the text, we refer to operations involving $l_{r}$ and $\tau$ as operations of a Learned Oracle. This is to abstract away from the exact details of the Learning Model when it is unnecessary for the results we are trying to prove. We define this term precisely here, to avoid confusion.
Definition 2.10 (Learned Oracle).
Let a Learned Oracle $L$ be an oracle that, given a Bloom Filter Learning Model $l_{r}$ and a threshold $\tau$, returns positive if $l_{r}(x)\geq\tau$. $L$ returns negative otherwise.
Definition 2.11 (Learned Construction).
We define $\tilde{C}_{r}$ to be a setup algorithm such that $\tilde{C}_{r}(1^{\lambda},S,\mathcal{D})=\tilde{M}_{r}^{S}$, where $\tilde{M}_{r}^{S}$ is a learned compressed representation of $S$ i.e a compressed representation that includes a Bloom Filter Learning Model.
Returning to our running example, consider the set $S=\{x,y\}$ and the dataset $\mathcal{D}=\{(x,1),(y,1),(z,0)\}$ in a Learned Bloom Filter that uses the Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ such that $l_{r}^{\mathcal{D}}(x)=0.6$ and $l_{r}^{\mathcal{D}}(y)=0.4$, and uses the threshold $\tau=0.5$. Let $M_{r}^{S^{\prime}}$ be the compressed representation of the set $S^{\prime}=\{x:x\in S|l(x)<\tau\}=\{y\}$ created by any setup algorithm for the Classical Bloom Filter. A trivial setup algorithm for the Learned Bloom Filter will then return the compressed representation $\tilde{M}_{r}^{S}=(l(x),\tau,M_{r}^{S^{\prime}})$.
Definition 2.12 (Learned Query).
We define $\tilde{Q}_{s}$ to be a query algorithm similar to $Q_{s}$ (Definition  2.2) differing only in that $\tilde{Q}_{s}$ only takes a learned compressed representation $\tilde{M}_{r}^{S}$ of the set $S$ instead of any compressed representation $M_{r}^{S}$.
In our running example, a trivial query algorithm, $\tilde{Q}_{s}(\tilde{M}_{r}^{S},x)$ would then be $l(x)\geq 0.5\lor(M_{r}^{S^{\prime}}[h_{1}(x)]=1\land M_{r}^{S^{\prime}}[h_{2}(x)]=1)$.
Definition 2.13 (Learned Bloom Filter).
Let a Learned Bloom Filter be a data structure $\tilde{B}=(\tilde{C}_{r},\tilde{Q_{s}})$ where $\tilde{C}_{r}$ obeys Definition 2.11 and $\tilde{Q}_{s}$ obeys Definition 2.12. The query algorithm $\tilde{Q}_{s}$ for the Learned Bloom Filter is
$\tilde{Q}_{s}=l(x)\geq\tau\lor Q_{s}(M_{r}^{S^{\prime}},x)=1$
where $Q_{s}$ is a query algorithm for the Classical Bloom Filter, and $M_{r}^{S^{\prime}}$ is the internal representation of the Backup Classical Bloom Filter encoding the set $S^{\prime}=\{x\in S|l_{r}(x)<\tau\}$
It is trivial to define a steady Learned Bloom Filter with completeness and soundness properties analogous to the ones outlined in Definition 2.5 for a Classical Bloom Filter. We now construct an adversarial model for the Learned Bloom Filter.
Challenge 3 (Learned Resilient).
We denote this challenge as ${\Lambda^{1l}}_{A,t}(\lambda)$.
  1. 1.
    $S\leftarrow A_{C}(1^{\lambda+n\log(u)})$
  2. 2.
    $\mathcal{D}\leftarrow\Delta_{r}(S)$
  3. 3.
    $\tilde{M}_{r}^{S}\leftarrow\tilde{C}_{r}(1^{\lambda+n\log(u)},S,\mathcal{D})$
  4. 4.
    $x^{*}\leftarrow A_{Q}^{\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S)$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)$.
  5. 5.
    If $x^{*}\notin S\cup\{x_{1},\ldots,x_{t}\}$ and $\tilde{Q}_{s}(M^{S}_{r},x^{*})=1$, output $1$. Otherwise, output 00.
Note that the adversary may not choose threshold $\tau$. If the adversary is allowed to choose threshold $\tau$, then Challenge 3 is easily succeeded by choosing $\tau=0$. We are now ready to formally define security for the steady Learned Bloom Filter.
Definition 2.14 (Learned Resilient).
Let an $(n,t,\epsilon)$-adversarial resilient steady Learned Bloom Filter be any steady Learned Bloom Filter for which it holds that, $\forall\lambda>n\in\mathbb{N},P[\Lambda^{1l}_{A,t}(\lambda)=1]\leq\epsilon$.
We now propose a stronger adversary that has access to the internal state of the Learned Bloom Filter.
Challenge 4 (Learned Reveal Resilient).
We denote this challenge as ${\Lambda^{2l}}_{A,t}(\lambda)$.
  1. 1.
    $S\leftarrow A_{C}(1^{\lambda+n\log(u)})$
  2. 2.
    $\mathcal{D}\leftarrow\Delta_{r}(S)$
  3. 3.
    $\tilde{M}_{r}^{S}\leftarrow\tilde{C}_{r}(1^{\lambda+n\log(u)},S,\mathcal{D})$
  4. 4.
    A is allowed access to $\tilde{M}_{r}^{S}$
  5. 5.
    $x^{*}\leftarrow A_{Q}^{\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)}(1^{\lambda+n\log(u)},S)$. $A_{Q}$ performs at most $t$ queries $x_{1},\ldots,x_{t}$ to $\tilde{Q}_{s}(\tilde{M}^{S}_{r},\cdot)$.
  6. 6.
    If $x^{*}\notin S\cup\{x_{1},\ldots,x_{t}\}$ and $\tilde{Q}_{s}(M^{S}_{r},x^{*})=1$, output $1$. Otherwise, output 00.
Definition 2.15 (Learned Reveal Resilient).
Let an $(n,t,\epsilon)$-adversarial reveal resilient steady Learned Bloom Filter be any steady Learned Bloom Filter for which it holds that, $\forall\lambda>n\in\mathbb{N},P[\Lambda^{2l}_{A,t}(\lambda)=1]\leq\epsilon$.
3 Secure Constructions for the Learned Bloom Filter
We attempt two solutions to the problem of securing the Learned Bloom Filter. Our first solution combines a Learned Bloom Filter technique called “sandwiching” introduced by Mitzenmacher [8] and the use of a pseudo-random permutation on the Bloom Filter input set first proposed by Naor and Yogev [3]. Our second solution relies on the (non Sandwiched) Learned Bloom Filter and instead employs the use of pseudo-random permutations to two Backup Classical Bloom Filters that we introduce as part of the construction.
In Section 3.1, we first provide a review of the setting from Section 2.2. Then we formally define a Sandwiched Learned Bloom Filter. Finally we provide a brief introduction to pseudo-random permutations. We introduce the Uptown Bodega Filter in Section 3.2. We then discuss concerns regarding the utility of the Uptown Bodega Filter which lead us to the Downtown Bodega Filter, which we introduce in Section 3.3.
3.1 Preliminaries
Setting Review: Consider a set of elements $\mathcal{K}\subset S$ and a set of elements $\mathcal{U}$ such that $\forall u\in\mathcal{U},u\notin S$. We form a dataset $\mathcal{D}=\{(x_{i},y_{i}=1)|x_{i}\in\mathcal{K}\}\cup\{(x_{i},y_{i}=0)|x_{i}\in\mathcal{U}\}$. We train a Bloom Filter Learning Model, $l_{r}^{\mathcal{D}}$, on the training dataset $\mathcal{D}$. Let $l_{r}^{\mathcal{D}}(x)$ be the probability estimate from the learning model (Definition 2.9) that $x$ is an element in $S$. A value $\tau$ may be chosen as a threshold. When $l_{r}^{\mathcal{D}}(x)\geq\tau$ then the Learned Bloom Filter considers $x$ to be an element of $S$. Otherwise, the Learned Bloom Filter passes $x$ on to the Backup Classical Bloom Filter.
Definition 3.1 (Sandwiched Learned).
Let a Sandwiched Learned Bloom Filter, $SB_{r}=(\tilde{C}_{r},\tilde{Q_{s}})$ be a data structure where $\tilde{C}_{r}$ obeys Definition 2.11, and $\tilde{Q}_{s}$ obeys Definition 2.12.
The learned compressed representation of any set $S$ and training dataset $\mathcal{D}$ under a Sandwiched Learned Bloom Filter consists of the following:
  1. 1.
    A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
  2. 2.
    A suitable threshold $\tau$ for $l_{r}^{\mathcal{D}}$
  3. 3.
    $M_{r}^{S}$, the compressed representation of the complete set $S$ encoded by a Classical Bloom Filter. We refer to this Classical Bloom Filter as the Initial Classical Bloom Filter.
  4. 4.
    $M_{r}^{S^{\prime}}$, the compressed representation of the set $S^{\prime}=\{x:x\in S|l_{r}^{\mathcal{D}}(x)<\tau\}$ encoded by a Classical Bloom Filter. We refer to this Classical Bloom Filter as the Backup Classical Bloom Filter.
The query algorithm $\tilde{Q}_{s}$ for the Sandwiched Learned Bloom Filter is
$(Q_{s}(M_{r}^{S},x)=1)\land(l_{r}^{\mathcal{D}}(x)>\tau\lor Q_{s}(M_{r}^{S^{\prime}},x)=1)$
where $Q_{s}$ is a query algorithm for the Classical Bloom Filter.
Figure 3: A Sandwiched Learned Bloom Filter. The initial filter only allows positives (true positive and false positive) to reach the Learned Bloom Filter.
Figure 3 shows an example of a Sandwiched Learned Bloom Filter. In our running example, the setup algorithm for a Sandwiched Learned Bloom Filter creates the same Bloom Filter Learning Model, $l_{r}$, as well as the same compressed representation for the Backup Classical Bloom Filter $M_{r}^{S^{\prime}}$ (trained on the set $S^{\prime}=\{y\}$) that we discussed after Definition 2.11. In addition, it also creates a representation $M_{r}^{S}$ for the Initial Classical Bloom Filter trained on the complete set $S=\{x,y\}$. A trivial query algorithm $\tilde{Q}_{s}(\tilde{M}_{r}^{S},x)$ for the Sandwiched Learned Bloom Filter would then be $(M_{r}^{S}[h_{1}(x)]=1\land M_{r}^{S}[h_{2}(x)]=1)\land(l_{r}(x)>\tau\lor(M_{r}^{S^{\prime}}[h_{1}(x)]=1\land M_{r}^{S^{\prime}}[h_{2}(x)]=1)$.
Lemma 3.2.
If $x$ is a false positive in a Sandwiched Learned Bloom Filter, $SB_{r}$, then $x$ is a false positive in the Initial Classical Bloom Filter of the Sandwiched Learned Bloom Filter.
Proof.
The proof follows from the definition of a Sandwiched Learned Bloom Filter. ∎
We use the standard definitions for pseudo-random permutations. We provide a brief but self-contained treatment adapted from Chapter 3 of Katz Lindell [13] here. Let $\text{Perm}_{n}$ be the set of all permutations on $\{0,1\}^{n}$.
Definition 3.3.
Let an efficient permutation $F$ be any permutation for which there exists a polynomial time algorithm to compute $F_{k}(x)$ given $k$ and $x$, and there also exists a polynomial time algorithm to compute $F_{k}^{-1}(x)$ given $k$ and $x$.
Definition 3.4.
Let $F:{\{0,1\}}^{*}\times{\{0,1\}}^{*}\mapsto{\{0,1\}}^{*}$ be an efficient, length-preserving, keyed function. $F$ is a keyed permutation if $\forall k$, $F_{k}(\cdot)$ is one-to-one.
Definition 3.5.
Let $F:{\{0,1\}}^{*}\times{\{0,1\}}^{*}\mapsto{\{0,1\}}^{*}$ be an efficient keyed permutation. $F$ is a pseudo-random permutation if for all probabilistic polynomial time distinguishers $D$, there exists a negiligible function $\mathtt{negl}$, such that
$|\Pr[D^{F_{k}(\cdot)F_{k}^{-1}(\cdot)}(1^{n})=1]-\Pr[D^{f_{n}(\cdot)f_{n}^{-1}(\cdot)}(1^{n})=1]|\leq\mathtt{negl}(n)$
where the first probability is taken over uniform choice of $k\in\{0,1\}^{n}$ and the randomness of $D$, and the second probability is taken over uniform choice of $f\in\text{Perm}_{n}$ and the randomness of $D$.
Theorem $4.8$ of Naor Yogev [3] proves that for a classical steady $(n,\epsilon)$-Bloom Filter that uses $m$ bits of memory, if pseudo-random permutations exist, then there exists a negiligible function $\mathtt{negl}$ such that for security parameter $\lambda$ there exists a $(n,\epsilon+\mathtt{negl}(\lambda))$-adversarial resilient Bloom Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory. This secure Bloom Filter can be constructed by running the initialization algorithm on $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ instead of $S$ [3].
3.2 The Uptown Bodega Filter
Now, we introduce our first secure construction for the Learned Bloom Filter built on top of the Sandwiched Learned Bloom Filter and a pseudo-random permutation. We call this construction an Uptown Bodega Filter.
Definition 3.6 (Uptown).
Let a Uptown Bodega Filter be a data structure $UB_{r,k}=(F_{k},\tilde{C}_{r},\tilde{Q_{s}})$ such that $F_{k}$ is a pseudo-random permutation and $(\tilde{C}_{r},\tilde{Q_{s}})$ is a Sandwiched Learned Bloom Filter representing the set $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ in the steady setting.
Theorem 3.7.
Let $SB_{r}$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, then there exists a negligible function $\mathrm{negl(\cdot)}$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial resilient Uptown Bodega Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory.
Our proof follows the same line of reasoning as Theorem $4.8$ of Naor Yogev [3] which proves that, if pseudo-random permutations exist, any probabilistic polynomial time adversary cannot distinguish between the Classical Bloom Filter with pseudo-random permutations and Challenge 1 by more than a negligible advantage. Due to Lemma 3.2, this result also holds for the Uptown Bodega Filter. The crux of the proof relies on the fact that running $F_{k}$ on an adversary’s queries permits us to consider the queries as random and not chosen adaptively by an adversary, while having no effect on the correctness of the Uptown Bodega Filter.
Proof.
We construct a Uptown Bodega Filter $UB_{k,r}$ from a Sandwiched Bloom Filter $SB_{r}$ as follows. Choose the first key $k\in{\{0,1\}}^{\lambda}$ from pseudo-random permutation, $F_{k}$, over ${\{0,1\}}^{\log|U|}$. Let $S_{k}=F_{k}(S)={\{F_{k}(x):x\in S\}}$. We initialize $SB_{R}$ with $S_{k}$. For each input $x$, our query algorithm outputs $SB_{k,r}(F_{k}(x))$. Just like in the classical Bloom Filter construction of [3], the only additional memory required is to store the key $k$ for $F_{k}$, which is $\lambda$ bits. Similarly, the running time of the query algorithm of the Uptown Bodega Filter is one pseudo-random permutation more than the running time of the query algorithm of the Sandwiched Learned Bloom Filter.
The completeness of $UB_{k,r}$ follows from the completeness of $SB_{k,r}$. If $x\in S$ then $SB_{k,r}$ was initialized with $F_{k}(x)$ which will return $1$ from the completeness of $SB_{r}$. The resilience of the construction follows from the following argument: Consider an experiment where $F_{K}$ in the Uptown Bodega Filter is replaced by a truly random permutation oracle $R(\cdot)$. Since $x$ has not been queried, we know that $R(x)$ is a truly random element that was not queried before, and we may think of it as chosen prior to the initialization of $SB_{k,r}$. From the soundness of $SB_{r}$ we get that the probability of $x$ being a false positive is at most $\epsilon$.
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Uptown Bodega Filter we constructed using $R(\cdot)$ and the Uptown Bodega Filter construction that uses the pseudo-random permutation $F_{k}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that $A$ can attack $UB_{k,r}$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run $A$ on $UB_{k,r}$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations. ∎
3.2.1 Utility of the Uptown Bodega Filter
While we have shown the existence of $(n,\epsilon)$-adversarial resilient Uptown Bodega Filters, it is unclear whether there is any utility in using an Uptown Bodega Filter over a secure Classical Bloom Filter. This is due to the fact that we train our learning model on $F_{k}(x)\in S$, a pseudo-random permutation of the set $S$, instead of training it directly on $S$. Intuitively, any structure in $S$ that can be “learned” by a Bloom Filter Learning Model is destroyed by taking the pseudo-random permutation of $S$. In other words, if pseudo-random permutations exist, then to any Bloom Filter Learning Model that runs in Probabilistic Polynomial Time (PPT), $F_{k}(x)\in S$ should appear the same as $|S|$ values sampled from a truly random oracle $R(\cdot)$. A precise proof for this is beyond the scope of this paper.
3.3 The Downtown Bodega Filter
To address the utility concerns of the Uptown Bodega Filter, we introduce a second Secure Learned Bloom Filter which we call the Downtown Bodega Filter.
Figure 4: A Downtown Bodega Filter. Both Positives and Negatives results for the Learned Oracle are routed to Backup Bloom Filters secured with Pseudorandom Permutations
Definition 3.8 (Downtown).
Let a Downtown Bodega Filter be a data structure
$DB_{r,k_{A},k_{B}}=(F_{k_{A}},F_{k_{B}},\tilde{C}_{r},\tilde{Q}_{s})$ where $\tilde{C}_{r}$ obeys Definition 2.11, $\tilde{Q}_{s}$ obeys Definition 2.12, and $F_{k_{A}},F_{k_{B}}$ are pseudo-random permutations. In other words, $(\tilde{C}_{r},\tilde{Q}_{s})$ form a Learned Bloom Filter.
The learned compressed representation of any set $S$ and training dataset $\mathcal{D}$ under an Downtown Bodega Filter consists of the following:
  1. 1.
    A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
  2. 2.
    A suitable threshold $\tau$ for $l$
  3. 3.
    $M_{r}^{S^{\prime\prime}}$, the compressed representation of the set $S^{\prime\prime}=\{F_{k_{A}}(x):x\in S|l_{r}^{\mathcal{D}}(x)\geq\tau\}$ encoded by a Classical Bloom Filter. We refer to this Classical Bloom Filter as Backup Classical Bloom Filter A.
  4. 4.
    $M_{r}^{S^{\prime}}$, the compressed representation of the set $S^{\prime}=\{F_{k_{B}}(x):x\in S|l_{r}^{\mathcal{D}}(x)<\tau\}$ encoded by a Classical Bloom Filter. We refer to this Classical Bloom Filter as Backup Classical Bloom Filter B.
The query algorithm $\tilde{Q}_{s}$ for the Downtown Bodega Filter is
$(l_{r}^{\mathcal{D}}(x)\geq\tau\land Q_{s}(M_{r}^{S^{\prime\prime}},F_{k_{A}}(x))=1)\lor(l_{r}^{\mathcal{D}}(x)<\tau\land Q_{s}(M_{r}^{S^{\prime}}(F_{k_{B}}(x))=1)$
Lemma 3.9.
Let $B_{r}$ be an $(n,\epsilon)$-Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, then there exists a negligible function $\mathrm{negl(\cdot)}$ such that for security parameter $\lambda$, Backup Classical Bloom Filter A and Backup Classical Bloom Filter B are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient Bloom Filters each using $m^{{}^{\prime}}=m+\lambda$ bits of memory.
Lemma 3.9 is just a rephrasing of Theorem $4.8$ of Naor Yogev [3] and follows directly from the theorem. We include a self-contained proof here for completeness. We also make it more evident that the proof holds not just for adversarial resilient Bloom Filters but also for adversarial reveal resilient Bloom Filters that provide guarantees under a strictly stronger adversarial model. It is important to note that, while the adversary has access to the internal representation of the Downtown Bodega Filter, the adversary does not have access to the secret keys $k_{A}$ and $k_{B}$. This is consistent with the formulation of Naor Yogev [3] who state at the end of Section 4 in their paper, with reference to Theorem $4.8$, “Notice that, in all the above constructions only the pseudo-random function (permutation) key must remain secret. That is, we get the same security even when the adversary gets the entire memory of the Bloom filter except for the PRF (PRP) key.”
Proof.
Let us consider Backup Classical Bloom Filter $A$ (a proof for Backup Bloom Filter $B$ can be constructed in exactly the same way). Our setup algorithm $\tilde{C}_{r}$ merely initializes Backup Classical Bloom Filter $A$ with $S^{\prime\prime}$. Our query algorithm on input $x$ queries for $x^{{}^{\prime}}=F_{k_{A}}(x)$. The only additional memory required is for storing $k_{A}$ which is $\lambda$ bits long.
The completeness follows from the completeness of the Classical Bloom Filter. The resilience of the construction follows from the following argument: consider an experiment where $F_{k_{A}}$ in Backup Classical Bloom Filter $A$ is replaced by a truly random oracle $\mathcal{R}(\cdot)$. Since $x$ has not been queried, we know that $R(x)$ is a truly random element that was not queried before, and we may think of it as chosen prior to the initialization of Backup Bloom Filter $A$. From the soundness of Backup Classical Bloom Filter $A$, we get that the probability of $x$ being a false positive is at most $\epsilon$.
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Backup Bloom Filter $A$ we constructed using $\mathcal{R}(\cdot)$ and the Backup Classical Bloom Filter $A$ construction that uses the pseudo-random permutation $F_{k_{A}}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that adversary $A$ can attack Backup Classical Bloom Filter $A$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run adversary $A$ on a Backup Classical Bloom Filter $A$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations. ∎
Theorem 3.10.
Let $SB_{r}$ be an $(n,\epsilon)$-Sandwiched Learned Bloom Filter using $m$ bits of memory. If pseudo-random permutations exist, then there exists a negligible function $\mathrm{negl(\cdot)}$ such that for security parameter $\lambda$ there exists an $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient Downtown Bodega Filter, $DB_{r,k_{A},k_{B}}$, that uses $m^{{}^{\prime}}=m+2\lambda$ bits of memory.
Proof.
We construct a Downtown Bodega Filter $DB_{r,k_{A},k_{B}}$ from a Sandwiched Learned Bloom Filter $SB_{r}$ as follows. We use the memory budget of the Initial Classical Bloom Filter to construct the Backup Classical Bloom Filter $A$. We use the memory budget of the Backup Classical Bloom Filter of the Sandwiched Learned Bloom Filter to construct Backup Classical Bloom Filter $B$. We do not modify the Bloom Filter Learning Model $l$ which remains trained on dataset $\mathcal{D}$ with threshold $\tau$. We choose keys $k_{A},k_{B}\in\{0,1\}^{\lambda}$ and use $2\lambda$ bits of extra memory to store them.
The completeness of $DB_{r,k_{A},k_{B}}$ follows from 1) the completeness of Backup Classical Bloom Filter $B$ and 2) the fact that any $x$ such that $l(x)<\tau$ is declared to be not in $S$ by $\tilde{Q}_{s}$ the query algorithm f the Downtown Bodega Filter if and only if the query algorithm of the Classical Bloom Filter with Backup Classical Bloom Filter $B$, $Q_{s}(M_{r}^{S^{\prime}},F_{k_{B}}(x))$, is also 00. This fact follows directly from the query algorithm $\tilde{Q}_{s}$ for the Downtown Bodega Filter stated in Definition 3.8.
To prove the resilience of the construction, we first show that the security of the Downtown Bodega Filter construction is reducible to the security of Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$. Consider a false positive i.e an $x\notin S$ for which the Downtown Bodega Filter returns $1$. From the definition of the Downtown Bodega Filter, one of the following two cases must be true.
Case 1: The Bloom Filter Learning Model, $l$, returned a value $\geq\tau$ and Backup Classical Bloom Filter $A$ returned $1$, more precisely, $l(x)\geq\tau\land Q_{s}(M_{r}^{S^{\prime\prime}},F_{k_{A}}(x))=1$
Case 2: The Bloom Filter Learning Model, $l$, returned a value $<\tau$ and Backup Bloom Classical Filter $B$ returned $1$, more precisely, $l(x)<\tau\land Q_{s}(M_{r}^{S^{\prime}},F_{k_{B}}(x))=1$
Therefore, for any probabilistic polynomial time (PPT) adversary to induce a false positive in the overall Downtown Bodega Filter construction, they must either induce a false positive in Backup Classical Bloom Filter $A$ or Backup Classical Bloom Filter $B$. We have already proven in Lemma 3.9 that both Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$ are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. It follows that the entire construction is $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. This concludes our proof. ∎
4 Discussion
In this section, we discuss the only two known attacks on the Learned Bloom Filter, introduced by Reviriego et al [2]. We refer to Attack 1 from their work as the Blackbox Mutation Attack, and Attack 2 from their work as the Whitebox Mutation Attack respectively. We discuss the Blackbox Mutation Attack in Section 4.1, and the Whitebox Mutation Attack in Section 4.2. Both sections include details on how our Secure Learned Bloom Filter constructions mitigate the attack.
4.1 Black-box Mutation Attack
The black-box adversarial model defined by Reviriego et al [2] is slightly weaker but very similar to the adversarial model we define in Challenge 3 of Section 2.2. Both our adversarial model and Reviriego et al’s black-box adversary model allows the adversary access to query the Learned Bloom Filter. One major difference is that our model allows the adversary to choose the initial set $S$ that is represented by the Learned Bloom Filter, whereas Reviriego et al’s model does not. In their attack, Reviriego et al first test elements until a positive (whether a false positive or true positive) is found. They then mutate the positive by changing a small fraction of the bits in the input in order to generate more false positives. The attack targets the Learned Oracle (recall Figure 2) by making it generate false positives without the input reaching the Backup Bloom Filter.
Since, as we mentioned, Reviriego et al’s attack relies on the fact that the adversary can query the Learned Oracle and make it generate false positives without any involvement of the Backup Bloom Filter, a Sandwiched Bloom Filter construction is intuitively more resilient against this kind of mutation attack. Since the Uptown Bodega Filter takes a pseudo-random permutation of the input $x$ prior to sending it to either the Learned Oracle or the Bloom Filters, the mutation approach is no longer viable. Unlike in the Uptown Bodega Filter, the Downtown Bodega Filter does give the adversary the ability to query the Learned Oracle directly. However, the extra (pseudo-random permuted) Backup Classical Bloom Filter for the case $l(x)>\tau$ lowers the probability of mutation-based false positives from the Learned Oracle ending up as false positives in the overall construction.
4.2 White-box Mutation Attack
The white-box adversarial model defined by Reviriego et al [2] is similar to the adversarial model we define in Challenge 4 in Section 2.2. With knowledge of the state of the Learned Oracle, the adversary can generate mutations in a more sophisticated way. Reviriego et al provide the example of a malicious URL dataset where an adversary may begin with a non-malicious URL and make changes such as removing the “s” in “https” or removing the “www” to generate false positives. Since we have shown the Downtown Bodega Filter to be $(n,t,\epsilon)$-adversarial reveal resilient in the steady setting, such mutations will not provide the adversary any advantage over the construction.
5 Hybrid Adversarial Model
In this section, we first define a hybrid model, where part of the queries are chosen by an adversary, while the rest are non-adversarial (“regular”) queries. Next we analyze the performance of the Downtown Bodega Filter and the Secure Classical Bloom Filter respectively in our hybrid model. We provide results for the conditions in which a Downtown Bodega Filter outperforms the Secure Classical Bloom Filter in the hybrid model. We then discuss the tradeoffs of our approach and provide a realistic example where the Downtown Bodega Filter displays better performance as compared to a Secure Classical Bloom Filter under many settings.
5.1 Hybrid Model
Let $A=(A_{C},A_{Q})$ be a probabilistic polynomial time (PPT) adversary as defined in Section 2. Consider a set of $N$ queries sent to a Bloom Filter. Our adversary, $A$, is allowed to choose exactly $\alpha N$ of those queries, where $\alpha\in[0,1]$. As an exampleto, the $N$ queries may be part of a streaming workload under any of the streaming models described by Muthukrishnan [14].
5.2 Downtown Bodega Filter
Figure 5: To generate a false positive in the Downtown Bodega Filter, the adversary must either 1) generate a false positive in the Learned Oracle and direct their query through Backup Bloom Filter $A$ or 2) generate a true negative in the Learned Oracle and direct their query through Backup Bloom Filter $B$
Let $DB_{r}=(\tilde{C}_{r},\tilde{Q}_{s})$ be a Downtown Bodega Filter. Let $l_{r}$ be the corresponding Bloom Filter Learning Model and $\tau$ its threshold. Let $M_{r}^{S^{\prime\prime}}$ be the internal representation of Backup Classical Bloom Filter $A$ and $M_{r}^{S^{\prime}}$ be the internal representation of Backup Classical Bloom Filter $B$. For an adversarial query to generate a false positive in the Downtown Bodega Filter, one of the following must hold true for a given query (see Figure 5):
  1. 1.
    The query generates a false positive in the Learned Oracle and a false positive in Backup Bloom Filter $A$.
  2. 2.
    The query generates a true negative in the Learned Oracle and a false positive in Backup Bloom Filter $B$
We first model the false positive probability of the Downtown Bodega Filter as a function of its memory budget, $m$, the set that the Downtown Bodega Filter is encoding, $S$, and the corresponding training dataset $\mathcal{D}$. Without loss of generality, let $\alpha_{P}$ of the adversarial queries be such that they generate false positives in the Learned Oracle and go through Backup Bloom Filter $A$. Similarly, let $\alpha_{N}$ of the adversarial queries be such that they generate true negatives and go through Backup Bloom Filter $B$. Note that $\alpha=\alpha_{P}+\alpha_{N}$.
Let $\text{FPR}(S,m)$ be the expected false positive probability of a Classical Bloom Filter that encodes the set $S$ with memory budget $m$. Let $\text{FPR}_{L}(S,\mathcal{D},m)$ be the expected false positive probability of a Learned Oracle $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. Similarly, let $\text{TNR}_{L}(S,\mathcal{D},m)$ be the expected true negative probability of a Learned Oracle $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. We make the assumption that the correctness probability of the Learned Oracle is independent of the correctness probability of the Backup Classical Bloom Filters. In particular, we assume that for any $m,m^{\prime}$, $\text{FPR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{FPR}_{L}(m)\text{FPR}(m^{\prime})$ and $\text{TNR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{TNR}_{L}(m)\text{FPR}(m^{\prime})$.
Consider a system where the total memory budget is $M$. Let the memory allocation of a Downtown Bodega Filter from memory budget $M$ be assigned as follows. Let $m_{L}$ be the number of bits of memory assigned to a Learned Oracle $L$. Let $m_{A}$ be the number of bits of memory assigned to Backup Bloom Filter $A$. Let $m_{B}$ be the number of bits of memory assigned to Backup Bloom Filter $B$. Let $\lambda$ be the number of bits assigned to the key of the pseudo-random permutations used prior to Backup Bloom Filter $A$ and Backup Bloom Filter $B$. Note that to stay within the memory budget it must hold that $M\geq m_{L}+m_{A}+m_{B}+2\lambda$.
Theorem 5.1.
For any memory budget $M$, any set $S$, and corresponding training dataset $\mathcal{D}$, The Downtown Bodega Filter encoding $S$ provides an expected false positive probability of
$\text{FPR}_{DB}(S,\mathcal{D},M)=\text{FPR}_{L}(S,\mathcal{D},m_{L})\text{FPR}(S^{\prime\prime},m_{A})+\text{TNR}_{L}(S,\mathcal{D},m_{L})\text{FPR}(S^{\prime},m_{B})$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof.
From the definition of a Downtown Bodega Filter (Definition 3.8, in particular look at the formulation for the query algorithm $\tilde{Q}_{s}$. Figure 5 is also helpful here) it follows that a false positive in the overall construction must either be a false positive in Backup Classical Bloom Filter $A$ or a false positive in Backup Classical Bloom Filter $B$. If the query is a false positive in Backup Classical Bloom Filter $A$, it must also be a false positive in the Learned Oracle. Alternatively, if the query is a false positive in Backup Classical Bloom Filter $B$, it must be a true negative in the Learned Oracle. The result follows. ∎
We now derive an expression for the false positive probability of the Downtown Bodega Filter in the hybrid adversarial setting. Recall that we are assuming that out of $N$ queries, the adversary makes $\alpha_{P}$ queries that generate a false positive in the Learned Oracle and $\alpha_{N}$ queries that generate a true negative in the Learned Oracle.
Theorem 5.2.
In the hybrid adversarial setting, the expected false positive probability of the Downtown Bodega Filter is
$\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof.
For each query $i$ among $N$ queries, one of the following cases holds.
Case 1: The query is not an adversary generated query. Therefore as established by Theorem 5.1, the False Positive Probability for the query is $\text{FPR}_{DB}(S,\mathcal{D},M)$. There are $(1-\alpha_{P}-\alpha_{N})N$ such queries.
Case 2: The query is an adversary generated query such that it generates a false positive in the Learned Oracle. Since the Learned Oracle generating a false positive and the Learned Oracle generating a true negative are mutually exclusive events, the False Positive Probability for the query is the False Positive Probability of Backup Classical Bloom Filter $A$ i.e $\text{FPR}(S^{\prime\prime},m_{A})$. There are $\alpha_{P}N$ such queries.
Case 3: The query is an adversary generated query such that it generates a true negative in the Learned Oracle. Following logic similar to case 2, we can derive the False Positive Probability of the query to be $\text{FPR}(S^{\prime},m_{B})$. There are $\alpha_{N}N$ such queries.
The expected false positive probability of $N$ queries is then $\frac{1}{N}(\alpha_{P}N\cdot\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}N\cdot\text{FPR}(S^{\prime\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})N\cdot\text{FPR}_{DB}(S,M))$. The statement of the theorem follows. ∎
5.3 Secure Classical Bloom Filter
An alternative construction is to simply use a well-tuned $(n,t,\epsilon)$-adversarial reveal resilient Classical Bloom Filter. We will refer to this construction in this section as the Secure Classical Bloom Filter without causing confusion. The expected false positive probability of the Secure Classical Bloom Filter encoding a set $S$ is, by definition, $\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$. The extra $\lambda$ is due to the fact that the Secure Classical Bloom Filter requires one less pseudo-random permutation than the Downtown Bodega Filter.
We now provide an expression that encapsulates all the cases in the hybrid adversarial setting where a Downtown Bodega Filter construction provides a lower false positive probability compared to a secure Classical Bloom Filter construction for the same memory budget.
Theorem 5.3.
For a given memory budget $M=m_{L}+m_{A}+m_{B}+2\lambda$, any set $S$ and corresponding training dataset $\mathcal{D}$, in the hybrid adversarial setting with given $\alpha=\alpha_{P}+\alpha_{N}$, the expected false positive probability of the Downtown Bodega Filter is lower than the expected false positive probability of the Secure Classical Bloom Filter if the following holds true:
$\alpha_{P}\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}\text{FPR}(S^{\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})\text{FPR}_{DB}(S,\mathcal{D},M)<\text{FPR}(S,m_{L}+m_{A}+m_{B}+\lambda)$
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Bloom Filters $A$ and $B$, the random coins used in the construction of the Secure Classical Bloom Filter, and the random coins used in the generation of the non-adversarial queries.
Proof.
The proof follows directly from the expression derived for the expected false positive probability of the Downtown Bodega Filter construction in Theorem 5.2 and the expression for the expected false positive probability of the Secure Classical Bloom Filter. ∎
5.4 Realistic Example
Mitzenmacher Broder [7] show that the false positive probability for a Classical Bloom Filter with $m$ bits encoding a set $S$, using $k(S,m)$ hash functions is
$\textrm{FPR}(S,m)=(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
In our analysis, we use the value of the number of hash functions $k(S,m)$ is always optimally chosen to be $k(S,m)=\ln 2\cdot(m/|S|)$ (this optimal value is also derived by Mitzenmacher Broder [7]). We further model the false positive rate of a Learned Oracle as being the same as the false positive rate of a Classical Bloom Filter encoding set $S$ but with a better false positive probability for the same memory budget. This is consistent with the assumptions made by prior work including Kraska et al [1] and Mitzenmacher [8].
$\textrm{FPR}_{L}(S,\mathcal{D},m)=c(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)}$
where $c\leq 1$.
We note that the true negative probability of a Learned Oracle is merely the probability of a negative entry (which is constant as we are assuming set $S$ is constant) is not marked as a false positive. Let $Q_{N}$ be the fraction of true negative non-adversarial queries. We have
$\textrm{TNR}_{L}(S,\mathcal{D},m)=(1-\textrm{FPR}_{L}(S,\mathcal{D},m))Q_{N}=(1-c(1-e^{-k(S,m)\cdot|S|/m})^{k(S,m)})Q_{N}$
To evaluate how much lower the false positive probability of a Learned Oracle needs to be for the Downtown Bodega Filter to perform better than the Secure Classical Bloom Filter, we may then use these derivations in Theorem 5.3.
We choose realistic values for our example from prior work on evaluating Learned Bloom Filters [1] on Google’s transparency report. We pick $2$ Megabytes as our memory budget, $m$, chosen from the range of values in Figure 10 of Kraska et al [1]. We choose the cardinality of the set we want to encode, $|S|$, as $1.7$ million based on the number of unique URLs in Google’s transparency report evaluated in Kraska et al [1]. With a memory budget of $2$ Megabytes, Kraska et al [1] demonstrate that a Learned Bloom Filter has $0.25$ of the False Positive Ratio of a Classical Bloom Filter, hence we use that as our value for $c$.
Parameter Explanation Value
M Memory Budget 2 Megabytes
$m_{L}$ Memory budget for Learned Oracle 1 Megabytes
$m_{A}$ Memory budget for Backup Bloom Filter $A$ $0.5$ Megabytes
$m_{B}$ Memory budget for Backup Bloom Filter $B$ $0.5$ Megabytes
n Cardinality of set to encode 1.7 Million
c FPR of Learned Bloom Filter / FPR of Classical Bloom Filter 0.25
$\lambda$ Number of bits in secret key 128 bits
Table 1: A summary of the chosen values for our realistic example of the performance tradeoffs of a Downtown Bodega Filter compared to a Secure Classical Filter in the hybrid adversarial case
We use $128$ bits as the size of our security parameter, $\lambda$. For the case of the Downtown Bodega Filter, we let the Learned Oracle take $1$ Megabytes, while dividing the remaining $1$ Megabytes equally between Backup Classical Bloom Filters $A$ and $B$. Backup Classical Bloom Filters $A$ and $B$ encode $S^{\prime\prime}$ and $S^{\prime}$ respectively, which are both subsets of the set $S$ (refer to Definition 3.8). Since the false positive probability is adversely impacted by a larger set cardinality when the memory budget and the number of hash functions are kept constant, we take $|S^{\prime\prime}|=|S^{\prime}|=|S|$ as a lower bound on the performance of the Downtown Bodega Filter (i.e an upper bound on its false positive rate).
We take $\alpha$ to be a variable ranging from 00 to $0.2$ equally divided between $\alpha_{P}$ and $\alpha_{N}$. Our chosen values are summarized in Table 1. Figure 6 shows the results of our calculations. As can be seen, when the adversary has access to less than a certain cutoff fraction of the workload, the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter for the same memory budget. The source code for our analysis can be found in an anonymously hosted Google Colab notebook [15].
Figure 6: The False Positive Rate of the Downtown Bodega Filter compared to the Secure Classical Bloom Filter in the hybrid adversarial setting where the adversary chooses between 00 and $0.2$ of the workload.
6 Open Problems
This work suggests the following three open problems.
Problem 1: Secure Constructions for the Learned Bloom Filter in the unsteady setting. While we have provided secure constructions for the Learned Bloom Filter, our results only hold for the steady setting, where the query algorithm $\tilde{Q}_{s}$ may not modify the internal representation of the Learned Bloom Filter. Naor Yogev [3] provide secure constructions for the Classical Bloom Filter in an unsteady settings. We leave the formulation of secure constructions for the Learned Bloom Filter in the unsteady setting as an open problem.
Problem 2: Secure Learned Bloom Filters under a stronger adversarial model. The adversarial model we consider for the Learned Bloom Filter is stronger than Naor Yogev [3] in that it allows an adversary access to the internal state of the Bloom Filter, but it is weaker than Clayton et al [11] and Filic et al [12] as it does not allow the adversary to insert entries in the Learned Bloom Filter. We leave the formulation of secure constructions under an adversarial model that allows insertion as an open problem.
Problem 3: Secure Learned Bloom Filters under computationally unbounded adversaries. In this work, we only consider probabilistic polynomial time (PPT) adversaries. An open problem is to prove or disprove any security guarantees for our constructions against computationally unbounded adversaries. Alternatively, an open problem is to provide secure constructions against computationally unbounded adversaries. Naor Yogev [3] discuss computationally unbounded adversaries in Section 5 of their work.
7 Acknowledgements
The authors would like to thank Hussain Tirmazi for initial feedback on the paper. We thank the anonymous reviewers for constructive feedback and insightful comments.